OpenStudy (anonymous):

A 3m long ladder is leaned on a wall so that the foot of the ladder is 1m from the base of the wall. How high up the wall does the ladder reach?

OpenStudy (anonymous):

 Do you know what you need to apply here?

OpenStudy (anonymous):

That's what I need to find out.. I have an idea. X² = 1m² + 3m² X² = 1m + 9m X = 10m X = √10m X = 3.1622776601683793319988935444327 X = 3.16 (2DP)

OpenStudy (anonymous):

@Joq

OpenStudy (anonymous):

You're in the right path. Let me help you a little bit.

OpenStudy (anonymous):

You already know that by using the Pythagorean Theorem we can solve the problem: \(\large{a^2= b^2 + c^2}\) \(a\) is the hypotenuse, and \(b\) and \(c\) are the shortest sides. In this case, the hypotenuse is the ladder, and we also have the length of one of the shortest sides (the length made by the bottom of the ladder with the wall), let call it \(b\). So we need to find \(c\): \(\large{a= 3\ m}\) \(\large{b= 1\ m}\) \(\large{c= \ ?\ m}\) \(\large{(3\ m)^2= (1\ m)^2 + c^2}\) Then I think you can solve for c.

OpenStudy (anonymous):

You just need to isolate \(c\) and do the same thing you do above, and you're done.

OpenStudy (anonymous):

Do you think you could explain this a bit better? I'm 14.

OpenStudy (anonymous):

@Joq

OpenStudy (anonymous):

Sure! Let me try.

OpenStudy (anonymous):

In our problem, \(a\) and \(b\) are the knowns, and we need to find \(c\). In order to find \(c\) we need to isolate it in one side of the equation, to say, in the right-hand side, because we know that in the left-hand side its value is going to show up (since there are no more unknowns). How do we do that? We have: \(\large{a^2= b^2 + c^2}\) \(\large{a^2-b^2= c^2}\) <- Subtract \(b^2\) from both sides. \(\large{\sqrt{a^2-b^2}= c}\) <- Take the inverse operation of an exponent to the 2nd power, which is a square root, in order to find the result of \(c\) and not \(c^2\). If we try to do this: \(a = \sqrt{b^2 + c^2}\) it's like trying to find the hypotenuse, but by looking at the drawing, we do know the hypotenuse, which is 3m: the ladder.

OpenStudy (anonymous):

So I'd work it out like this: X² = 3m² - 1m² X² = 9m - 1m X = 8m X = √8m X = 2.8284271247461900976033774484194 X = 2 (NWN) @Joq

OpenStudy (anonymous):

Yep!. Exactly.

OpenStudy (anonymous):

I think the answer can be rounded to 2 decimal places (\(2.83\ m \)) or written with the square root (\(2\sqrt{2}\ m)\).

OpenStudy (anonymous):

Thank-you ever so much. That really helped a lot. Thanks :-)

OpenStudy (anonymous):

@Joq

OpenStudy (anonymous):

You're very welcome @ellieearnshaw (: